OpenStudy (anonymous):

what is the prime factor of 343

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (immanuelv):

7^3

OpenStudy (anonymous):

how can that be when 12 time 7 = 84

OpenStudy (immanuelv):

7*7*7 = 343

OpenStudy (immanuelv):

where did you get 12 from?

OpenStudy (immanuelv):

hope this helps!

OpenStudy (anonymous):

you should start dividing the number by a prime number, then divide THAT product by another prime number, and so on and so forth. For example: 343/7 = 49 49/7 = 7 Therefore the prime factors of 343 are 7*7*7. Got it? :)